ZEKE SPECTROSCOPY OF ALUMINUM AND GALLIUM COMPLEXES OF PIPERIDINE

\underline{PARAGRANJITA BHOWMIK}, SHENGGANG LI, BRADFORD R. SOHNLEIN, JASON F. FULLER, XU WANG, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506.

The Al and Ga complexes were produced by reactions of laser-vaporized metal atoms with piperidine (C_{5}H_{11}N) and characterized by pulsed field ionization-zero electron kinetic energy (PFI-ZEKE) spectroscopy. For Al-C_{5}H_{11}N, the adiabatic ionization potential was measured to be 37060 cm^{−1}, and vibrational frequencies were determined for the Al^{+}- N stretching (238 cm^{−1}), Al^{+}/Al-N-C bending (84/105 cm^{−1}) and piperidine-based vibrations (300 and 464 cm^{−1}). By comparing the ZEKE measurements with ab initio calculations, the neutral and ionic aluminum complexes were determined to be in C_{s} symmetry with Al binding to the nitrogen of the chair-like piperidine ring. For the gallium complex, the ZEKE spectrum shows an ionization potential of 37529 cm^{−1}, a Ga^{+}-N stretching frequency of 176 cm^{−1}and a Ga^{+}-N-C bending frequency of 70 cm^{−1}.